# CoreMark: Toward Robust and Universal Text Watermarking Technique

### v1

@@ V-C Resistance to Potential Adaptive Attacks
Resistance to Key-Recovery. In general, the watermark information is first encrypted before being embedded into the host medium [7, 12]. In CoreMark, we employ BLAKE2 [41] as a cryptographic hash function, using both the secret key ($k$) and the watermark as inputs to generate an encrypted watermark. Since attackers lack knowledge of the specific parameters used, the key $k$ remains undisclosed. The probability of successfully breaking the key is $2^{\text{-len(key)}}$, indicating that longer keys provide stronger security assurances.
@@ V-D Ablation Study
Effects of Quantization Step $t$. In our main results, the quantization step size $t$ is set to 1. Since varying $t$ values correspond to different embedding strengths, we further investigate their impact on CoreMark’s performance. Specifically, we evaluate the robustness (e.g., screenshot, print-scan) and imperceptibility of the method under three embedding strengths: $t=1$, $2$, and $3$, corresponding to Step 1, 2, 3, respectively. The experiments are conducted using the Calibri font across a range of font sizes to examine the performance variations in both aspects. As illustrated in Figure 14, increasing the embedding strength consistently enhances robustness but slightly reduces visual quality. Nevertheless, CoreMark achieves a good balance between robustness and imperceptibility. Users can flexibly select the embedding strength based on application requirements: for scenarios where visual quality is critical, Step 1 or Step 2 yields a PSNR above 33 dB and SSIM exceeding 0.998 across most font sizes; for more adversarial environments, Step 3 offers near-perfect extraction accuracy (approaching 100%). By default, $t$ is set to 1. We also analyze the effect of the parameter $T_{c}$ in Figure 15, finding that while it has little impact on invisibility, overly small values may weaken robustness. Thus, setting a relatively higher value for $T_{c}$ is sufficient to ensure both invisibility and robustness.
@@
Effects of ES Modulator. To objectively assess the core performance of CoreMark, the ES modulator is excluded from all main experiments except Table III. To further verify the generality of ES modulator mentioned in Section IV-A2, we apply it to both our CoreMark and Tan19 [8], and evaluate its impact on robustness across different font sizes. As shown in Figure 16, for Tan19, the addition of the ES modulator significantly improves the extraction accuracy, especially for smaller font sizes, with the average accuracy increasing from approximately 65% to over 95%. For CoreMark, which already demonstrates strong robustness, the ES modulator further enhances performance, maintaining extraction accuracy close to 100% across all font sizes. These results highlight the effectiveness and generalizability of the ES modulator in enhancing watermark robustness under challenging conditions.
@@ VII Conclusion
In this paper, we analyzed existing text watermarking methods and found that they struggle to balance invisibility and robustness, with occasional embedding failure. We also revealed that methods relying on edge features are particularly vulnerable to binarization. Based on our analysis, we propose a new embedding paradigm termed core to balance robustness and invisibility while maintaining generalizability across diverse languages and fonts. Building upon this paradigm, we present CoreMark, a text watermarking framework encompassing both embedding and extraction processes. The embedding process first extracts cores from individual characters and selects characters with superior robustness, then embeds watermarks by thickening or thinning cores. The extraction process is achieved by comparing core widths with an adaptive threshold. Furthermore, we propose a general plug-and-play embedding strength modulator to enhance robustness for small-sized characters by adaptively adjusting the embedding strength. Experimental results show that CoreMark achieves stronger robustness with better invisibility compared with state-of-the-art methods, while maintaining satisfying generalizability.

### v2

@@ V-C Resistance to Potential Adaptive Attacks
Resistance to Key-Recovery. In general, the watermark information is first encrypted before being embedded into the host medium [7, 12]. In CoreMark, we employ BLAKE2 [41] as a cryptographic hash function, using both the secret key and the watermark as inputs to generate an encrypted watermark. Since attackers lack knowledge of the specific parameters used, the key remains undisclosed. The probability of successfully breaking the key is $2^{\text{-len(key)}}$, indicating that longer keys provide stronger security assurances.
@@ V-D Ablation Study
Effects of Quantization Step $\beta$. In our main results, the quantization step size $\beta$ is set to 1. Since varying $\beta$ values correspond to different embedding strengths, we further investigate their impact on CoreMark’s performance. Specifically, we evaluate the robustness (e.g., screenshot, print-scan) and imperceptibility of the method under three embedding strengths: $\beta=1$, $2$, and $3$, corresponding to Step 1, 2, 3, respectively. The experiments are conducted using the Calibri font across a range of font sizes to examine the performance variations in both aspects. As illustrated in Figure 14, increasing the embedding strength consistently enhances robustness but slightly reduces visual quality. Nevertheless, CoreMark achieves a good balance between robustness and imperceptibility. Users can flexibly select the embedding strength based on application requirements: for scenarios where visual quality is critical, Step 1 or Step 2 yields a PSNR above 33 dB and SSIM exceeding 0.998 across most font sizes; for more adversarial environments, Step 3 offers near-perfect extraction accuracy (approaching 100%). By default, $\beta$ is set to 1. We also analyze the effect of the parameter $T_{c}$ in Figure 15, finding that while it has little impact on invisibility, overly small values may weaken robustness. Thus, setting a relatively higher value for $T_{c}$ is sufficient to ensure both invisibility and robustness.
@@
Effects of ES Modulator. To objectively assess the core performance of CoreMark, the ES modulator is excluded from all main experiments except Table IV. To further verify the generality of ES modulator mentioned in Section IV-A2, we apply it to both our CoreMark and Tan19 [8], and evaluate its impact on robustness across different font sizes. As shown in Figure 16, for Tan19, the addition of the ES modulator significantly improves the extraction accuracy, especially for smaller font sizes, with the average accuracy increasing from approximately 65% to over 95%. For CoreMark, which already demonstrates strong robustness, the ES modulator further enhances performance, maintaining extraction accuracy close to 100% across all font sizes. These results highlight the effectiveness and generalizability of the ES modulator in enhancing watermark robustness under challenging conditions.
@@ VII Conclusion
In this paper, we analyzed existing text watermarking methods and found that they struggle to balance invisibility and robustness, with occasional embedding failure. We also revealed that methods relying on edge features are particularly vulnerable to binarization. Based on our analysis, we propose a new embedding paradigm termed core to balance robustness and invisibility while maintaining generalizability across diverse languages and fonts. Building upon this paradigm, we present CoreMark, a text watermarking framework encompassing both embedding and extraction processes. The embedding process first extracts cores from individual characters and selects characters with superior robustness, then embeds watermarks by reducing or expanding cores. The extraction process is achieved by comparing core widths with an adaptive threshold. Furthermore, we propose a general plug-and-play embedding strength modulator to enhance robustness for small-sized characters by adaptively adjusting the embedding strength. Experimental results show that CoreMark achieves stronger robustness with better invisibility compared with state-of-the-art methods, while maintaining satisfying generalizability.